# Measuring kinetic inductance and superfluid stiffness of two-dimensional superconductors using high-quality transmission-line resonators

### v2

@@ IV.2 Kinetic inductance of a niobium diselenide flake
Fig. 5 displays optical micrographs of an exfoliated NbSe2 strip connected to a transmission line resonator. The fabrication starts with mechanical exfoliation of NbSe2 flakes onto SiO2/Si substrates in an inert N2 environment inside a glovebox with oxygen levels lower than 0.5 ppm. We employ the dry pickup technique Wang et al. (2013) using polycarbonate stamps to transfer NbSe2 flakes from the SiO2/Si to the destination area on a substrate with pre-patterned Nb resonator structures. After removing polycarbonate residue with chloroform, we pattern the connection between the NbSe2 flake and Nb transmission lines by electron-beam lithography. Next, we perform in situ Ar ion milling to remove surface oxide layers from the Nb resonator and NbSe2 flake (Fig. 5B), followed by subsequent deposition of a 3-nm-thick Ti sticking layer and reactive sputtering of 200–240 nm of NbN to make galvanic connection between the two with negligible contact resistanceAntony et al. (2022). We use a relatively thick NbN not only to ensure that it forms contact over the $\sim$100–300 nm thick NbSe2, but also to reduce the contribution of kinetic inductance from NbN by suppressing its sheet resistance.
@@
Fig. 8 plots the frequency of the NbSe2 resonator vs. temperature, with the contribution from the transmission line subtracted. The coherence length of NbSe2 is only roughly 10 nm, and therefore sets the shortest length scale between coherence length ($\zeta$) and penetration depth of the NbSe2 sample flake. Consequently, the superconducting NbSe2 is within the local limit ($\zeta\ll\lambda_{L}$), and we apply Eqn. 20 in order to infer $\mathcal{D}_{s}$ and $\lambda_{L}$. As shown in the right $y$-axis in Fig. 8A, $\mathcal{D}_{s}/k_{B}$ remains larger than the $T_{c}$ of NbSe2. At the zero-temperature limit, $\lambda_{L}\approx$ 440 nm, which is consistent with the reported valuesSonier et al. (1997); Fletcher et al. (2007). Any discrepancy between our reported value of penetration depth and previously reported values is probably due to the limitation of using Eqn. 19 to describe the two-gap NbSe2 superconductor, as well as error in the frequency subtraction.
@@
Fig. 8B plots $Q_{\rm i}$ of both the sample component and reference. Unlike in the temperature range plotted in Fig. 4B, the $Q_{\rm i}$ of the reference resonators drops considerably because we are now closer to the $T_{c}$ of Nb. In this temperature range, $Q_{\rm i}$ of the sample component and reference are approximately the same, suggesting that $Q_{\rm i}$ is dominated by the $Q_{\rm 0}$ term rather than the $Q_{\text{eff}}$ in Eqn. 10. Since the participation ratio of the sample is merely a few ppt, the probe resonator quality factor $Q_{\rm 0}$ is too low for us to discern the true $Q$ factor of the sample component. This substantiates the limitation of this measurement technique: $Q_{\rm i}$ is only representative of the loss through the sample component when $Q_{\rm 0}\gg Q_{\rm eff}$ in the perturbation regime in which the participation ratio of the sample component is small.
@@ VII.1 Theory of kinetic inductance and its relation to the penetration depth
where $K_{\mu\nu}({\bf r}-{\bf r}^{\prime})$ is the Mattis-Bardeen kernel; this object is otherwise known as the current-current correlator, and can be computed via standard means using the Kubo formula. By matching the electromagnetic energy stored to that of the inductor, we get a general formula of the kinetic inductance
@@
where $\mathcal{D}_{s}$ is the superfluid stiffness. This quantity is calculated by computing the momentum dependent quantity $K_{\mu\nu}(\bf q)$ and replacing it with its $q\rightarrow 0$ limit. This expression is valid in the limit where the coherence length $\xi$ is much smaller than the London penetration depth; when $\xi$ is large, the momentum dependence of $K_{\mu\nu}(\bf q)$ can no longer be neglected. As for the vector potential, we can revert Eqn. 23 to write the kinetic inductance in terms of current density and then make the substitution
@@
The idea of penetration volume is extremely useful in understanding anisotropic responses Prozorov et al. (2000). Anisotropic superconductors have a direction dependent superfluid stiffness, that is, Eqn. 25 is modified to
@@ VII.2.1 Quarter-wave resonator
The $Q$ factor of the probe superconducting resonator, $Q_{0}$, is given by Eqn. 6.
To quantify the kinetic inductance and loss of the sample component using a quarter-wave resonator, we compare Eqn. 7 and 34 to find the frequency shift and change of $Q$ factor by identifying:
@@ VII.2.2 Half-wave resonator
The $Q$ factor of the bare probe resonator is given by Eqn. 12.
Comparing Eqn. 13 with 38,

### v3

@@ IV.2 Kinetic inductance of a niobium diselenide flake
Fig. 5 displays optical micrographs of an exfoliated NbSe2 strip connected to a transmission line resonator. The fabrication starts with mechanical exfoliation of NbSe2 flakes onto SiO2/Si substrates in an inert N2 environment inside a glovebox with oxygen levels lower than 0.5 ppm. We employ the dry pickup technique Wang et al. (2013) using polycarbonate stamps to transfer NbSe2 flakes from the SiO2/Si to the destination area on a substrate with pre-patterned Nb resonator structures. After removing polycarbonate residue with chloroform, we pattern the connection between the NbSe2 flake and Nb transmission lines by electron-beam lithography. Next, we perform in situ Ar ion milling to remove surface oxide layers from the Nb resonator and NbSe2 flake (Fig. 5B), followed by subsequent deposition of a 3-nm-thick Ti sticking layer and reactive sputtering of 200–240 nm of NbN to make galvanic connection between the two with negligible contact resistance Antony et al. (2022). We use a relatively thick NbN not only to ensure that it forms contact over the $\sim$100–300 nm thick NbSe2, but also to reduce the contribution of kinetic inductance from NbN by suppressing its sheet resistance. In addition to the resonator shown in Fig. 5, each NbSe2 device also consists of a Nb reference resonator and an Nb-NbN control resonator (see Fig. 17 in Appendix). No aluminum is used for the NbSe2 devices.
@@
Fig. 8 plots the frequency of the NbSe2 resonator vs. temperature, with the contribution from the transmission line subtracted. The coherence length of NbSe2 is only roughly 10 nm, and therefore sets the shortest length scale between coherence length ($\zeta$) and penetration depth of the NbSe2 sample flake. Consequently, the superconducting NbSe2 is within the local limit ($\zeta\ll\lambda_{L}$), and we apply Eqn. 19 in order to infer $\mathcal{D}_{s}$ and $\lambda_{L}$. As shown in the right $y$-axis in Fig. 8A, $\mathcal{D}_{s}/k_{B}$ remains larger than the $T_{c}$ of NbSe2. At the zero-temperature limit, $\lambda_{L}\approx$ 440 nm, which is consistent with the reported values Sonier et al. (1997); Fletcher et al. (2007). Any discrepancy between our reported value of penetration depth and previously reported values is probably due to the limitation of using Eqn. 18 to describe the two-gap NbSe2 superconductor, as well as error in the frequency subtraction.
@@
Fig. 8B plots $Q_{\rm i}$ of both the sample component and reference. Unlike in the temperature range plotted in Fig. 4B, the $Q_{\rm i}$ of the reference resonators drops considerably because we are now closer to the $T_{c}$ of Nb. In this temperature range, $Q_{\rm i}$ of the sample component and reference are approximately the same, suggesting that $Q_{\rm i}$ is dominated by the $Q_{\rm 0}$ term rather than the $Q_{\text{eff}}$ in Eqn. 9. Since the participation ratio of the sample is merely a few ppt, the probe resonator quality factor $Q_{\rm 0}$ is too low for us to discern the true $Q$ factor of the sample component. This substantiates the limitation of this measurement technique: $Q_{\rm i}$ is only representative of the loss through the sample component when $Q_{\rm 0}\gg Q_{\rm eff}$ in the perturbation regime in which the participation ratio of the sample component is small.
@@ VII.1 Theory of kinetic inductance and its relation to the penetration depth
where $K_{\mu\nu}({\bf r}-{\bf r}^{\prime})$ is the Mattis-Bardeen kernel Mattis and Bardeen (1958); this object is otherwise known as the current-current correlator, and can be computed via standard means using the Kubo formula. The response is generically non-local – i.e., the current at position $\bf r$ depends on the vector potential at $\bf{r}\neq\bf{r}^{\prime}$. The origin of non-locality is that the response to an electromagnetic field of a superconductor is spread over a region of size determined by the coherence length $\xi$. In terms of Fourier components, this non-locality brings in a variation in $K({\bf q})$ at the scale of $\xi$.
By matching the electromagnetic energy stored to that of the inductor, we get a general formula of the kinetic inductance
@@
where $\mathcal{D}_{s}$ is the superfluid stiffness. This quantity is calculated by computing the momentum dependent quantity $K_{\mu\nu}(\bf q)$ and replacing it with its $q\rightarrow 0$ limit. This expression is valid in the limit where the coherence length $\xi$ is much smaller than the London penetration depth; when $\xi$ is large, the momentum dependence of $K_{\mu\nu}(\bf q)$ can no longer be neglected. As for the vector potential, we can revert Eqn. 22 to write the kinetic inductance in terms of current density and then make the substitution
@@
The idea of penetration volume is extremely useful in understanding anisotropic responses Prozorov et al. (2000). Anisotropic superconductors have a direction dependent superfluid stiffness, that is, Eqn. 24 is modified to
@@ VII.2.1 Quarter-wave resonator
The $Q$ factor of the probe superconducting resonator, $Q_{0}$, is given by Eqn. 5.
To quantify the kinetic inductance and loss of the sample component using a quarter-wave resonator, we compare Eqn. 6 and 33 to find the frequency shift and change of $Q$ factor by identifying:
@@ VII.2.2 Half-wave resonator
The $Q$ factor of the bare probe resonator is given by Eqn. 11.
Comparing Eqn. 12 with 37,
@@ VII.3 Supplementary information
  6. 6.
    Frequency and quality factor temperature dependence data for three separate NbSe2 samples.
  7. 7.
    Optical image of NbSe2 device including sample, reference, and control resonators.
@@
Figure 16: Frequency and quality factor data for three different NbSe2 samples.
Figure 17: Optical image of NbSe2 device made from sputtered niobium and etched into three difference resonators: (A) reference, (B) sample, and (C) control. (A) is composed of niobium and silicon, (B) is a niobium resonator connected to an exfoliated flake of NbSe2 via Ar ion-milling followed by niobium nitride sputtering, and (C) is a niobium resonator with niobium nitride stitching used in (B), with no exfoliated flake.